# Modelling Sampling Distributions of Test Statistics with Autograd

### v1

@@ Abstract
Simulation-based inference methods that feature correct conditional coverage of confidence sets based on observations that have been compressed to a scalar test statistic require accurate modelling of either the p-value function or the cumulative distribution function (cdf) of the test statistic. If the model of the cdf, which is typically a deep neural network, is a function of the test statistic then the derivative of the neural network with respect to the test statistic furnishes an approximation of the sampling distribution of the test statistic. We explore whether this approach to modelling conditional 1-dimensional sampling distributions is a viable alternative to the probability density-ratio method, also known as the likelihood-ratio trick. Relatively simple, yet effective, neural network models are used whose predictive uncertainty is quantified through a variety of methods.
@@ 1 Introduction
Automatic differentiation (see, for example, Ref.[1]) has revolutionized machine learning, permitting the routine application of gradient descent algorithms to fit to data models of essentially unlimited complexity. The same technology can be used to take the derivative of these models with respect to their inputs without the need to explicitly calculate the derivatives [2]. A potentially useful application of this capability is approximating the probability density function (pdf), $f(x\mid\boldsymbol{\theta})$, given an accurate neural network model of the associated conditional cumulative distribution function (cdf), $F(x\mid\boldsymbol{\theta})$, using the fact that
@@
where $\boldsymbol{\theta}$ are the parameters of the data-generation mechanism, which we distinguish from the parameters $\mathbf{w}$ of the neural network model. This paper explores this possibility in the context of simulation-based frequentist inference [3, 4, 5, 6, 7].
Equation (1) furnishes an approximation of the pdf $f(x\mid\boldsymbol{\theta})$ whether $x$ is a function of the underlying observations $D$ only or if $x=\lambda(D;\boldsymbol{\theta})$ is a test statistic that depends on $D$ as well as on the parameters $\boldsymbol{\theta}$. Moreover, computing the derivative of the cdf using autograd to obtain the pdf is exact; autograd does not use finite difference approximations.
@@
Given the pdf $f(x\mid\boldsymbol{\theta})$, where the data $D$ have been compressed to the 1-dimensional quantity $x$, further inferences can be performed by treating the pdf as a statistical model for $x$. These include constructing approximate confidence intervals by profiling  [8] $f(x\mid\boldsymbol{\theta})$ or constructing posterior densities for the parameters $f(\boldsymbol{\theta}\mid x)$ given a suitable prior. However, if $x=\lambda$ is a test statistic some care must be exercised because, in general, a variation of $\boldsymbol{\theta}$ will induce a change in $\lambda$, which should be taken into account. The key issue is whether a sufficiently accurate model of the cdf $F(x\mid\boldsymbol{\theta})$ can be constructed. This is the issue we explore in this paper where the focus is on test statistics that arise in simulation-based inference. We consider the classic ON/OFF problem of astronomy  [9] and high-energy physics as a benchmark example and apply the insights gained to the SIR problem [10, 11] in epidemiology.
This paper is structured as follows. In Sec. 2 we review related work in the areas of smiulation-based inference and uncertainty quantification in neural networks. In Sec. 3 we model the pdf from the cdf of a test statistic for problem where the likelihood is tractable: the prototypical signal/background problem in high energy physics, which in astronomy is known as the On/Off problem. In sec. 4 we model the pdf from the cdf of a test statistic for a problem where the likelihood is not tractable: the SIR model in epidemiology. In both examples, various techniques are applied to improve the accuracy and quantify the uncertainty. The paper ends with a discussion and conclusion in Sec. 5.
@@ 2.1 Likelihood Ratio Trick
It has been known since at least the early 1990s [12, 13, 14, 15] that probabilities can be modeled with neural networks using the density ratio method, also known as the likelihood ratio trick [3, 16, 17, 18, 19]. Suppose one has two data samples, $\{(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})\}_{A}$ and $\{(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})\}_{B}$. Sample $A$ comprises pairs $(\boldsymbol{x}_{i},\boldsymbol{\theta}_{i})$ in which the components are sampled sequentially: $\boldsymbol{\theta}_{i}\sim\pi$ followed by $\boldsymbol{x}_{i}\sim G(\boldsymbol{\theta}_{i})$, where $\pi$ is a known prior and $G$ is a simulator. Sample $B$ differs from $A$ in that $\boldsymbol{x}_{i}\sim g$, where $g(\boldsymbol{x}\mid\boldsymbol{\theta})$ is a known density that may, or may not, depend on $\boldsymbol{\theta}$. If sample $A$ is assigned target $t=1$ and $B$ is assigned $t=0$, then a sufficiently flexible function — fitted by minimizing the cross-entropy loss averaged over a large balanced training sample — yields an approximation to the function
$D(\boldsymbol{x},\boldsymbol{\theta})$ $=\frac{f(\boldsymbol{x},\boldsymbol{\theta})}{f(\boldsymbol{x},\boldsymbol{\theta})+g(\boldsymbol{x}\mid\boldsymbol{\theta})\pi(\boldsymbol{\theta})},$ (2)
where $f(\boldsymbol{x},\boldsymbol{\theta})$ is the (generally unknown) joint density associated with sample $A$. A rearrangement of Eq.(2) leads to the result
$f(\boldsymbol{x}\mid\boldsymbol{\theta})$ $=\frac{f(\boldsymbol{x},\boldsymbol{\theta})}{\pi(\boldsymbol{\theta})},$
$=g(\boldsymbol{x}\mid\boldsymbol{\theta})\left(\frac{D}{1-D}\right).$ (3)
@@ 2.2 LF2I and ALFFI
Our motivation for starting with the cdf (or p-value) is that constructing a model of it is the key step in the likelihood-free frequentist (LF2I) approach  [6] and in an extension of it (ALFFI)  [7] in which the test statistic $x=\lambda(D;\boldsymbol{\theta})$ is included as an input to the neural network-based model of the cdf. If the test statistic $\lambda$ is chosen so that large values of $\lambda$ disfavor the hypothesis $H:\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$ it follows that a confidence set, $R(D)$, can be constructed at confidence level $\tau=1-\alpha$ where $\alpha$ is the miscoverage rate. By definition the set $R(D)$ is all values of $\boldsymbol{\theta}_{0}$ for which $F(\lambda\mid\boldsymbol{\theta}_{0})\leq\tau$ given data $D$ [6, 7]. The fact that $\lambda$ is an input to the model in the ALFFI algorithm presents an opportunity: taking the derivative of the approximate cdf $\hat{F}$ with respect to $\lambda$ provides an approximation $\hat{f}(\lambda\mid\boldsymbol{\theta})$ of the sampling distribution of the test statistic. In this paper we explore the accuracy with which both the cdf and pdf can be approximated.
@@ 2.3 Conformal Inference
2.4 Multistage Modelling of Neural Networks
@@
2.5 Bootstrap NNs
The bootstrap [29] is a statistical method that treats an observed dataset as if it were a population. This makes it possible, for example, to approximate the sampling distributions of statistics. Given an estimator $\hat{\theta}$ of $\theta$ the bootstrap quantifies the uncertainty in $\hat{\theta}$ by repeatedly sampling the original dataset with replacement to create new datasets. The uncertainty in $\hat{\theta}$ can be quantified with a measure of the variability of predictions between different bootstrap datasets [30, 31].
@@ 3 Example 1: signal/background or ON/OFF model
where $\mu$ and $\nu$ are the mean signal and background counts, respectively, and we use the same likelihood ratio test statistic as in [7]
@@
3.1 Modelling the CDF with ALFFI
ALFFI is an algorithm for approximating the conditional cdf, $F(x\mid\theta)$, of a scalar random variable $x$ where $\theta$ denotes the parameters of the statistical model. If $x=\lambda(D;\theta)$ is a test statistic then the cdf can be used to construct confidence sets for all parameters simultaneously. Here we explore the accuracy of the ALFFI algorithm for modelling the cdf with a view to deriving the sampling distribution, $f(\lambda\mid\mu,\nu)$ of the ON/OFF test statistic by differentiating the cdf with respect to $\lambda$.
Again following [7] we start by sampling $\mu$ and $\nu$ from uniform priors and sample $n_{i}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $m_{i}\sim\text{Poiss}(\nu_{i})$. At each parameter point, we calculate $\lambda_{i}=\lambda(n_{i},m_{i}\mid\mu_{i},\nu_{i})$ according to Eq. 10 and the procedure is repeated to sample $N_{i},M_{i}$ and $\lambda_{D}=\lambda(N_{i},M_{i}\mid\mu_{i},\nu_{i})$. Finally, the indicator $Z$, which is unity if $\lambda_{i}\leq\lambda_{D}$ and zero otherwise, is computed. This results in a training set of size $B$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},Z_{i})\}_{i=1}^{B}$. The observation in ALFFI and LF2I is that critical value functions, such as the cdf $F(\lambda\mid\mu,\nu)=\mathbb{P}(\lambda\leq\lambda_{D}\mid\mu,\nu)$, are the expectation value $\mathbb{E}(Z\mid\lambda_{D},\mu,\nu)$ of the discrete random variable $Z$, of which a smooth approximation can be created with a deep neural network $f(\mathbf{x}_{i},\mathbf{w})$. Critically, this network must be trained to minimize the mean square error (MSE) loss [^1] In LF2I and ALFFI, the predicted value is $\mathbb{E}[Z\mid x]=\mathbb{P}(Z=1\mid x)=\frac{\mathbb{P}(x\mid Z=1)\mathbb{P}(Z=1)}{\mathbb{P}(x\mid Z=1)+\mathbb{P}(x\mid Z=0)}$. In other words, it behaves like a classifier, and because of that the cross entropy loss can also be used.,
@@ 3.1 Modelling the CDF with ALFFI
The cdf, using the ALFFI algorithm, was approximated using a fully connected neural network with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$, 12 hidden layers with 12 nodes each, and a single output. A sigmoid was used in the output layer to constrain the output to the unit interval. The activation function at each hidden node is a SiLU [32]. The SiLU nonlinearity was used in all the models in this study, since the SiLU allows the model to be smooth and differentiable multiple times. The network was trained with the NAdam optimizer [33] with a fixed learning rate of $6\times 10^{-4}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $512$ for $10^{5}$ iterations, that is, for 5 epochs. Longer training runs were also performed, but did not yield improvements in the results.
@@
Figure 1: ON/OFF cdfs modelled with ALFFI at different $(\mu,\nu)$ points.
@@
Figure 2: ON/off pdfs obtained by differentiating cdfs modelled with ALFFI at different $(\mu,\nu)$ points.
@@
Although the ALFFI algorithm yields valid confidence sets for the ON/OFF problem, it does not yield a sufficiently accurate approximation of the cdf, and hence of the pdf. This is shown in Fig. 1, where the empirical cdf [^2] Suppose that we have a sample of observations $\lambda_{1},...,\lambda_{n}$. The empirical cdf is defined by $F(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(\lambda_{i}\leq\lambda\right)$, $F$, is displayed together with the approximated cdf $\hat{F}_{\text{ALFFI}}(\lambda\mid\mu,\nu)$. Taking the derivative of the approximated cdf yields the ALFFI approximation of the pdf, $\hat{f}_{\text{ALFFI}}(\lambda)=\frac{\partial\hat{F}_{\text{ALFFI}}}{\partial\lambda}$, displayed in Fig. 2. We observe that $\hat{f}_{\text{ALFFI}}(\lambda)$ displays sharp fluctuations and turning points, which was also confirmed by taking a numerical derivative of $\hat{F}_{\text{ALFFI}}$. This is because the actual slope of the cdf changes dramatically from point to nearby point, especially at low values of $\lambda$. It can also be seen that the histogrammed pmf of $\lambda$ at a particular $(\mu,\nu)$ point is a high frequency function, or one that is noisy and features very sharp peaks. The fluctuations seen in the pmf also depend on the choice of binning: finer binning shows larger fluctuations because this is a discrete distribution. This high-frequency nature of the pmf complicates our attempt to create a smooth modelling of the pmf with ALFFI.
3.2 Directly modelling the empirical cdf

### v2

# Modeling Sampling Distributions of Test Statistics with Autograd

@@ Abstract
Simulation-based inference methods that feature correct conditional coverage of confidence sets based on observations that have been compressed to a scalar test statistic require accurate modeling of either the p-value function or the cumulative distribution function (cdf) of the test statistic. If the model of the cdf, which is typically a deep neural network, is a function of the test statistic then the derivative of the neural network with respect to the test statistic furnishes an approximation of the sampling distribution of the test statistic. We explore whether this approach to modeling conditional 1-dimensional sampling distributions is a viable alternative to the probability density-ratio method, also known as the likelihood-ratio trick. Relatively simple, yet effective, neural network models are used whose predictive uncertainty is quantified through a variety of methods.
@@ 1 Introduction
Automatic differentiation (see, for example, Ref.[1]) has revolutionized machine learning, permitting the routine application of gradient descent algorithms to fit to data models of essentially unlimited complexity. The same technology can be used to take the derivative of these models with respect to their inputs without the need to explicitly calculate the derivatives, [2] which can be cumbersome or intractable for complex models. In this paper we leverage this capability to investigate whether it is possible to obtain an accurate approximation of the probability density function (pdf), $f(x\mid\boldsymbol{\theta})$, given a neural network model of the associated conditional cumulative distribution function (cdf), $F(x\mid\boldsymbol{\theta})$, using the fact that
@@
where $\boldsymbol{\theta}$ are the parameters of the data-generation mechanism, which we distinguish from the parameters $\mathbf{w}$ of the neural network (NN) model. A pdf $f(x\mid\boldsymbol{\theta})$ is the key ingredient in both frequentist and Bayesian inference, while the cdf arises naturally in the context of simulation-based frequentist inference [3, 4, 5, 6, 7]. Simulation-based inference (SBI) makes it possible to perform inferences without the need for explicit knowledge of the statistical model. However, given that SBI often furnishes the cdf $F(x\mid\boldsymbol{\theta})$, standard inferential methods may be brought to bare if one had access to the associated pdf. For example, it would be possible to construct confidence intervals for any parameter of the model or deploy methods such as maximum likelihood.
Equation (1) furnishes an approximation of the pdf $f(x\mid\boldsymbol{\theta})$ whether $x$ is a function of the underlying observations $D$ only or if $x=\lambda(D;\boldsymbol{\theta})$ is a test statistic that depends on $D$ as well as on the parameters $\boldsymbol{\theta}$. Moreover, computing the derivative of the cdf using automatic differentiation (or autograd) to obtain the pdf is exact; autograd does not use finite difference approximations. Our main contributions are as follows.
  • We suggest a novel use of automatic differentiation.
  • We investigate the feasibility of accurately approximating the pdf by leveraging a neural network model of the cdf.
  • We explore multiple techniques for improving the accuracy of both the cdf and pdf, and methods for uncertainty quantification including the use of conformal inference.
  • We provide insights into modeling challenges.
@@
Given the pdf $f(x\mid\boldsymbol{\theta})$, where the data $D$ have been compressed to the 1-dimensional quantity $x$, further inferences can be performed by treating the pdf as a statistical model for $x$. These include constructing approximate confidence intervals by profiling  [8] $f(x\mid\boldsymbol{\theta})$ or constructing posterior densities for the parameters $f(\boldsymbol{\theta}\mid x)$ given a suitable prior. However, if $x=\lambda$ is a test statistic some care must be exercised because, in general, a variation of $\boldsymbol{\theta}$ will induce a change in $\lambda$, which should be taken into account. Since the quality of the estimated pdf is inherently tied to the accuracy of the cdf model, the key issue is whether a sufficiently accurate model of the cdf $F(x\mid\boldsymbol{\theta})$ can be constructed. This is the issue we explore in this paper where the focus is on test statistics that arise in simulation-based inference. We consider the classic ON/OFF problem of astronomy  [9] and high-energy physics as a benchmark example and apply the insights gained to the SIR problem [10, 11] in epidemiology.
This paper is structured as follows. In Sec. 2 we review related work in the areas of smiulation-based inference and uncertainty quantification in neural networks. In Sec. 3 we model the pdf from the cdf of a test statistic for a problem where the likelihood is tractable: the prototypical signal/background problem in high-energy physics, which in astronomy is known as the ON/OFF problem. In sec. 4 we model the pdf from the cdf of a test statistic for a problem where the likelihood is intractable: the SIR model in epidemiology. In both examples, various techniques are applied to improve the accuracy of the modeled pdf and quantify the uncertainty. The paper ends with a discussion and conclusion in Sec. 5.
@@ 2.1 Likelihood Ratio Trick
It has been known since at least the early 1990s [12, 13, 14, 15] that probabilities can be modeled with neural networks using the density ratio method, also known as the likelihood ratio trick [3, 16, 17, 18, 19]. Suppose one has two data samples, $\{(\mathbf{x}_{i},\boldsymbol{\theta}_{i})\}_{A}$ and $\{(\mathbf{x}_{i},\boldsymbol{\theta}_{i})\}_{B}$. Sample $A$ comprises pairs $(\mathbf{x}_{i},\boldsymbol{\theta}_{i})$ in which the components are sampled sequentially: $\boldsymbol{\theta}_{i}\sim\pi$ followed by $\mathbf{x}_{i}\sim G(\boldsymbol{\theta}_{i})$, where $\pi$ is a known prior and $G$ is a simulator. Sample $B$ differs from $A$ in that $\mathbf{x}_{i}\sim g$, where $g(\mathbf{x}\mid\boldsymbol{\theta})$ is a known density that may, or may not, depend on $\boldsymbol{\theta}$. If sample $A$ is assigned target $t=1$ and $B$ is assigned $t=0$, then a sufficiently flexible function — fitted by minimizing the cross-entropy loss averaged over a large balanced training sample — yields an approximation to the function
$D(\mathbf{x},\boldsymbol{\theta})$ $=\frac{f(\mathbf{x},\boldsymbol{\theta})}{f(\mathbf{x},\boldsymbol{\theta})+g(\mathbf{x}\mid\boldsymbol{\theta})\pi(\boldsymbol{\theta})},$ (2)
where $f(\mathbf{x},\boldsymbol{\theta})$ is the (generally unknown) joint density associated with sample $A$. A rearrangement of Eq.(2) leads to the result
$f(\mathbf{x}\mid\boldsymbol{\theta})$ $=\frac{f(\mathbf{x},\boldsymbol{\theta})}{\pi(\boldsymbol{\theta})},$
$=g(\mathbf{x}\mid\boldsymbol{\theta})\left(\frac{D}{1-D}\right).$ (3)
@@
2.2 Flow-based Methods
Many methods based on normalizing flows estimate a conditional pdf, for example, see [20]. Normalizing flows exploit the change of variables formula
$f(\mathbf{x}\mid\boldsymbol{\theta})$ $=g(z=h(\mathbf{x})|\boldsymbol{\theta})|\det J|,$ (4)
by modeling the function $h:X\in\mathbb{R}^{d}\rightarrow Z\in\mathbb{R}^{d}$ with a sequence of bijections, where $g(*)$ is a tractable pdf and $|\det J|$ is the determinant of the Jacobian, $J$, of the transformation.
@@ 2.3 LF2I and ALFFI
Our motivation for starting with the cdf (or p-value) is that constructing a model of it is the key step in the likelihood-free frequentist (LF2I) approach  [6] and in an extension of it (ALFFI)  [7] in which the test statistic $x=\lambda(D;\boldsymbol{\theta})$ is included as an input to the neural network-based model of the cdf. If the test statistic $\lambda$ is chosen so that large values of $\lambda$ disfavor the hypothesis $H:\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$ it follows that a confidence set, $R(D)$, can be constructed at confidence level $\tau=1-\alpha$ where $\alpha$ is the miscoverage rate. By definition, the set $R(D)$ is all values of $\boldsymbol{\theta}_{0}$ for which $F(\lambda\mid\boldsymbol{\theta}_{0})\leq\tau$ given data $D$ [6, 7]. The fact that $\lambda$ is an input to the model in the ALFFI algorithm presents an opportunity: taking the derivative of the approximate cdf $\hat{F}$ with respect to $\lambda$ provides an approximation $\hat{f}(\lambda\mid\boldsymbol{\theta})$ of the sampling distribution of the test statistic. In this paper we explore the accuracy with which both the cdf and pdf can be approximated.
@@ 2.4 Conformal Inference
2.5 Multistage modeling of Neural Networks
@@
2.6 Bootstrap Neural Networks
The bootstrap [30] is a statistical method that treats an observed dataset as if it were a population. This makes it possible, for example, to approximate the sampling distributions of statistics. Given an estimator $\hat{\theta}$ of $\theta$, the bootstrap quantifies the uncertainty in $\hat{\theta}$ by repeatedly sampling the original dataset with replacement to create new datasets. The uncertainty in $\hat{\theta}$ can be quantified with a measure of the variability of predictions between different bootstrap datasets [31, 32].
@@ 3 Example 1: signal/background or ON/OFF model
where $\mu$ and $\nu$ are the mean signal and background counts, respectively. The same likelihood ratio test statistic is used as in [7]
@@
3.1 Modeling the cdf with ALFFI
ALFFI is an algorithm for approximating the conditional cdf, $F(x\mid\theta)$, of a scalar random variable $x$ where $\theta$ denotes the parameters of the statistical model. If $x=\lambda(D;\theta)$ is a test statistic then the cdf can be used to construct confidence sets for all parameters simultaneously. Here we explore the accuracy of the ALFFI algorithm for modeling the cdf with a view to deriving the sampling distribution, $f(\lambda\mid\mu,\nu)$ of the ON/OFF test statistic by differentiating the cdf with respect to $\lambda$.
Again following [7] we start by sampling $\mu$ and $\nu$ from uniform priors and sample $n_{i}\sim\text{Poiss}(\mu_{i}+\nu_{i})$ and $m_{i}\sim\text{Poiss}(\nu_{i})$. At each parameter point, we calculate $\lambda_{i}=\lambda(n_{i},m_{i}\mid\mu_{i},\nu_{i})$ according to Eq. 11 and the procedure is repeated to sample $N_{i},M_{i}$ and $\lambda_{D}=\lambda(N_{i},M_{i}\mid\mu_{i},\nu_{i})$. Finally, the indicator $Z$, which is unity if $\lambda_{i}\leq\lambda_{D}$ and zero otherwise, is computed. This results in a training set of size $B$, $\mathcal{T}=\{(\mu_{i},\nu_{i},\lambda_{i},Z_{i})\}_{i=1}^{B}$. The key observation in ALFFI and LF2I is that critical value functions, such as the cdf $F(\lambda\mid\mu,\nu)=\mathbb{P}(\lambda\leq\lambda_{D}\mid\mu,\nu)$, are the expectation value $\mathbb{E}(Z\mid\lambda_{D},\mu,\nu)$ of the discrete random variable $Z$, of which a smooth approximation can be created with a deep neural network $f(\mathbf{x}_{i},\mathbf{w})$. Critically, this network must be trained to minimize the mean square error (MSE) loss [^1] In LF2I and ALFFI, the predicted value is $\mathbb{E}[Z\mid x]=\mathbb{P}(Z=1\mid x)=\frac{\mathbb{P}(x\mid Z=1)\mathbb{P}(Z=1)}{\mathbb{P}(x\mid Z=1)+\mathbb{P}(x\mid Z=0)}$. In other words, it behaves like a classifier, and because of that the cross entropy loss can also be used.,
@@ 3.1 Modeling the cdf with ALFFI
The cdf, using the ALFFI algorithm, was approximated using a fully-connected neural network with 3 input features $\mathbf{x}=\{\mu,\nu,\lambda\}$, 12 hidden layers with 12 nodes each, and a single output. A sigmoid was used in the output layer to constrain the output to the unit interval. The activation function at each hidden node is a SiLU [33]. The SiLU nonlinearity was used in all the models in this example, since the SiLU allows the model to be smooth and differentiable multiple times. The network was trained with the NAdam optimizer [34] with a fixed learning rate of $6\times 10^{-4}$. The training set is composed of $10^{7}$ examples, which were used in batches of size $512$ for $10^{5}$ iterations, that is, for 5 epochs. Longer training runs were also performed, but did not yield improvements in the results.
@@
Figure 1: ON/OFF problem: cdfs modeled with ALFFI at different $(\mu,\nu)$ points.
@@
Figure 2: ON/OFF problem: pdfs obtained by differentiating cdfs modeled with ALFFI at different $(\mu,\nu)$ points.
@@
Although the ALFFI algorithm yields valid confidence sets for the ON/OFF problem, it does not yield a sufficiently accurate approximation of the cdf, and hence of the pdf. This is shown in Fig. 1, where the empirical cdf [^2] Suppose that we have a sample of observations $\lambda_{1},...,\lambda_{n}$. The empirical cdf is defined by $F(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(\lambda_{i}\leq\lambda\right)$, $F$, is displayed together with the approximated cdf $\hat{F}_{\text{ALFFI}}(\lambda\mid\mu,\nu)$. Taking the derivative of the approximated cdf yields the ALFFI approximation of the pdf, $\hat{f}_{\text{ALFFI}}(\lambda)=\frac{\partial\hat{F}_{\text{ALFFI}}}{\partial\lambda}$, displayed in Fig. 2. We observe that $\hat{f}_{\text{ALFFI}}(\lambda)$ displays sharp fluctuations and turning points, which was also confirmed by taking a numerical derivative of $\hat{F}_{\text{ALFFI}}$. This is because the actual slope of the cdf changes dramatically from point to nearby point, especially at low values of $\lambda$. It can also be seen that the histogrammed pmf of $\lambda$ at a particular $(\mu,\nu)$ point is a high frequency function, or one that is noisy and features very sharp peaks. The fluctuations seen in the pmf also depend on the choice of binning: finer binning shows larger fluctuations because this is a discrete distribution. This high-frequency nature of the pmf complicates our attempt to create a smooth modeling of the pmf with ALFFI.
3.2 Directly Modeling the Empirical CDF